a rock is thrown straight upward and it rises to a height of 20m. with what speed was it thrown



1
Expert's answer
2021-03-17T16:53:09-0400

"h=\frac{v^2-v_0^2}{2g}=\frac{0^2-v_0^2}{-2g}=\frac{v_0^2}{2g}\to"


"v_0=\sqrt{2gh}=\sqrt{2\cdot9.8\cdot 20}=19.8\ (m/s)" . Answer